OpenStudy (anonymous):

use sum of a geometric series to find the sum of the first 10 terms of this sequence an=27(0.1)^(n-1)

OpenStudy (anonymous):

30

OpenStudy (anonymous):

\[27\times \frac{1-.1^{10}}{1-.1}\]

OpenStudy (anonymous):

How did you use sum of a geometric series @PROSS

OpenStudy (anonymous):

which is fact is almost exactly 30

OpenStudy (anonymous):

ignore the \[.1^{10}=.0000000001\] and you get \[27\times \frac{1}{.9}=27\times \frac{1}{\frac{9}{10}}=27\times \frac{10}{9}=30\]

OpenStudy (anonymous):

use the formula saellite73 gives for an analytic approach. You can check your answer with a calculator.... sum(seq(Y1,x,1,10)) Where Y1=27(0.1)^(x-1)